OpenStudy (anonymous):

geometry the triangle has an area of 40 cm2 find heigh n of the triangle number theorm.find two consecutive even integers whose product is 168

OpenStudy (ganpat):

is this a single question ??

OpenStudy (ganpat):

A = 1/2 * base * height.. and, two consecutive integers.. x and (x+2) so x (x+2) = 168... x2 + 2x = 168 x2 + 2x - 168 = 0 x2 + 14x - 12x - 168 = 0 x (x + 14) -12 (x + 14) = 0 (x + 14) ( x - 12) = 0 so, x = 12.. So the two numbers are , 12 and 14...

OpenStudy (anonymous):

THANK YOU

OpenStudy (ganpat):

well, m not sure on the 1st part of question..

OpenStudy (anonymous):

THATS OK

OpenStudy (ganpat):

:) WC

OpenStudy (anonymous):

:)